OpenStudy (anonymous):

Which equation would best help solve the following problem? Maria releases a javelin 1.5 meters above the ground with an initial vertical velocity of 22 meters per second. How long will it take the javelin to hit the ground? –4.9t2 + 22t + 1.5 = 0 –4.9t2+ 22t = 1.5 –4.9t2 + 1.5t + 22 = 0 –16t2 + 22t + 1.5 = 0

OpenStudy (anonymous):

Y = -4.9 t^2 + 22t + 1.5 = 0. Find t Solve the quadratic equation for t. D = b^2 - 4ac = 513.4 --> VD = 22.66 t = 22/9.8 + 22.66/9.8 = 4.56 sec.